Todd has a $900 monthly buget and wants to decrease it by 25%, he has already decreased it 15%, what is his monthly savings going to be?
1
Expert's answer
2012-01-06T08:07:33-0500
Todd has a $900 monthly buget and wants to decrease it by 25%, he has already decreased it 15%, what is his monthly savings going to be?

Savings after decreasing 15%: 900*0.15;
rests: 900-900*0.15;
savings after decreasing 25% more: (900-900*0.15)*0.25 + 900*0.15 = 900*(1-0.15)*0.25 = 900*0.85*0.25 + 900*0.15 = 191.25 + 135 = 326.25.